Find the seventh term of an A.P whose first term is 102 and common difference is -3
1
Expert's answer
2020-06-08T17:46:56-0400

"a_1= 102; d= -3"

"a_7=a_1+6d = 102-18=84"

Thus, seventh term of the AP is 84.